# Seeing Straight: Document Orientation Detection for Efficient OCR

### v1

@@ A.1 Model Architecture
We adapt the image encoder from the Phi-3.5 model abdin2024phi3 as the backbone for our rotation classification task. To adapt the encoder for classification, we append a lightweight two-layer classification head consisting of a linear projection, followed by a GELU activation, dropout, and a final linear layer. This classification head operates on a pooled image representation and outputs logits over four rotation classes: -90°, 0°, 90°and 180°.
@@
Prior to encoding, input images undergo a dynamic cropping strategy inspired by the InternVL2-style sampling method internvl2_github. Specifically, each image is first scaled and padded to the closest HD resolution such that both height and width are divisible by 336. The padded image is then divided into non-overlapping crops of size 336×336, with a maximum of 16 crops per image. In addition to these local crops, we include a globally resized version of the full image (also 336×336) to capture holistic context. Each of these crops, including the resized full image, is independently passed through the image encoder, yielding 577 tokens (based on $14\times 14$ patch sizes) per crop, including the [CLS] token, each with a hidden dimension of 1024. We extract the [CLS] token from each encoded crop and compute their mean across all crops and the full image. This averaged [CLS] representation is then used as input to the classification head.
A.2 Implementation Details
@@
We performed all experiments in PyTorch (paszke2019pytorch) using HuggingFace Transformers (wolf2019huggingface) library (v4.41.2). Training employs the AdamW (kingma2020method) optimizer with a constant learning rate of $1\times 10^{-5}$, and cross-entropy loss. We used a batch size of 256 for variants without dynamic cropping, and 64 for those with dynamic cropping. To improve stability, gradient clipping with a maximum norm of 1.0 was applied. Each model is trained for up to 35 epochs with early stopping based on validation accuracy. We use a lightweight classification head (two fully connected layers with non-linear activations) to predict one of four rotation classes. All training was performed on a single NVIDIA H100 GPU, which has an average latency of 0.5s per image during inference. The same model has a latency of around 0.9s on a single NVIDIA A100 GPU. We used the implementation of OCRBenchV2777https://github.com/Yuliang-Liu/MultimodalOCR/blob/main/OCRBench_v2/eval_scripts/vqa_metric.py to compute the ANLS metrics. We adapted this script to report both word-level and character-level ANLS metrics. For structured analysis on ORB-SROIE, we test the model against 4 fields: company, date, address and total for the percentage match calculations. We provide more details about the configurations used across different baselines.
Tesseract smith2007overview
@@
LLaMA-4 (Maverick) 17B touvron2024llama4
@@
Gemini-2.5 Flash google2025gemini2.5flash
For the Gemini-2.5 Flash baseline, we used Google’s multimodal model accessed via the google.generativeai Python SDK. The model gemini-pro-vision was configured with an authenticated API key and executed in streaming mode for OCR text extraction. Each image was read as binary data and sent to the model as an inline MIME-typed input. The outputs were returned as plain-text responses and stored incrementally in JSONL format for further evaluation.
@@
  • Access: google.generativeai SDK
@@
VLM Rotation Classifier prompt: Classify this image into 1 of 4 classes of rotation. If the image of the document is rotated anticlockwise by 90 degrees, classify it as -90. If it is rotated by 90 degrees clockwise, classify it as 90. If it is upright and not rotated, classify it as 0, and if it is rotated by 180 degrees, classify it as 180. Reply with just the class number (-90, 0, 90, or 180) without any additional text. English and Multilingual OCR prompt: You are an OCR engine. You job is to extract the exact text from the given image. Preserve the text formatting, including line breaks, spaces, and any symbols or special characters. Respond in plain text format.
A.3 Data Sources and Augmentation
For English training data, we use documents from various categories such as invoices, rent agreements, electricity bills, and gas bills, collected from internal repositories. For Indic data, we curate document images from 11 languages (Hindi, Bengali, Tamil, Malayalam, Telugu, Marathi, Kannada, Gujarati, Punjabi, Oriya, and Assamese) using Wikisource. To ensure fair evaluation, we strictly exclude all test images from the training sets. Specifically, the ORB-En and ORB-Indic benchmarks are fully disjoint from the training data. Moreover, we do not use the SROIE 2019 training split during model training. All results reported are therefore in a zero-shot setting, highlighting the generalization capability of the proposed method across both English and Indic scripts in our benchmark.
@@ A.4 Evaluation and Baselines
    ORB-En: A 874-image benchmark composed of 374 images from the SROIE 2019 test set and 500 images from the SynthDog validation set, uniformly rotated and re-annotated.
@@
    ORB-Indic: A 328-image benchmark covering 11 Indic scripts, each manually annotated and verified for rotation labels, along with corresponding ground-truth transcriptions for OCR evaluation.
@@
  • Tesseract smith2007overview
  • TrOCR li2021trocr
  • docTR liao2023doctr
  • Nanonets-OCR-s nanonets2024ocr
  • Gemma-3 27B team2024gemma
  • LLaMA-4 (Maverick) 17B touvron2024llama4
  • GPT-4o hurst2024gpt4o
@@
  • Gemini-2.5 Pro google2025gemini2.5pro
  • Gemini-2.5 Flash google2025gemini2.5flash
A.5 Pipeline Integration
@@
To measure downstream impact, we insert our rotation classifier before OCR engines. We compare OCR related metrics with and without rotation correction. Field-level accuracy is used for structured forms like ORB-En (SROIE), and string alignment metrics are used for the free-text Indic dataset ORB-Indic and the ORB-En (SynthDog) data.
Appendix B Ablation Study
We also performed an ablation study and analyze three variants of our model in Table 6:
  1. 1.
  2. 2.
  3. 3.
Model Variant ORB-En (%) ORB-Indic (%)
Single-layer classifier (w/ dynamic cropping) 90.66 82.01
Multi-layer classifier (w/o dynamic cropping) 92.65 86.59
Multi-layer classifier with dynamic cropping (Ours) 96.81 92.68
Table 6: Ablation results comparing classifier complexity and dynamic cropping on ORB-English and ORB-Indic.
Our ablation study highlights the individual contributions of dynamic cropping and classifier complexity to overall model performance. The baseline variant with a single-layer classifier and dynamic cropping achieves moderate accuracy, indicating that while spatial diversity in input patches is beneficial, a shallow classification head may under utilize the rich representations from the Phi-3.5 encoder. Replacing it with a deeper multi-layer classifier improves results across both ORB-English and ORB-Indic datasets, even without cropping, suggesting that classifier depth significantly enhances discriminative capacity. Finally, combining both techniques i.e., dynamic cropping and a multi-layer classifier achieves the highest accuracy on both benchmarks (96.81% on ORB-En and 92.68% on ORB-Indic), demonstrating that dynamic cropping provides complementary spatial context that further boosts performance when paired with a sufficiently expressive classifier. This underscores the importance of architectural synergy in maximizing the rotation classification capability of vision-language models.
Appendix C Ethics Statement
This work centers on the responsible creation of a benchmark for evaluating OCR robustness to rotation across English and Indic languages. All data used in this study originates from publicly available sources such as Wikisource and established OCR datasets (e.g., SROIE, SynthDog). We ensured that no sensitive or personally identifiable information (PII) was collected or included. The Indic subset, derived from Wikisource, contains only publicly released literary and historical documents that are already digitized and verified. As the research involves only publicly available and non-personal data, formal IRB approval was not required. All external datasets, libraries, and tools are appropriately credited through citations. We exercised care to ensure fairness across scripts and languages, though residual bias may persist due to unequal digital representation among Indic languages. The resulting OCR-Rotation-Bench (ORB) aims to support the development of robust, multilingual, and rotation-aware OCR systems for broader real-world document understanding.
Appendix D Reproducibility Statement
@@
Table 7: Performance comparison on ORB-En (SynthDog) and ORB-Indic documents with and without the proposed rotation correction pipeline, alongside ideal conditions. We report Word Error Rate (WER) and Character Error Rate (CER) on free-form outputs from both open-source and closed-source models (lower is better), computed using the HuggingFace evaluate function. Second best scores are underlined.
@@
Table 8: Performance comparison on ORB-En (SynthDog) and ORB-Indic documents with and without the proposed rotation correction pipeline, alongside ideal conditions. We report percentage match scores at word and character level (higher is better). Second best scores are underlined.

### v2

@@ 10.1 Model Architecture
We adapt the image encoder from the Phi-3.5 model [1] as the backbone for our rotation classification task. To adapt the encoder for classification, we append a lightweight two-layer classification head consisting of a linear projection, followed by a GELU activation, dropout, and a final linear layer. This classification head operates on a pooled image representation and outputs logits over 12 rotation classes: 0°, 30°, 60°and so on, till 330°.
@@
Prior to encoding, input images undergo a dynamic cropping strategy inspired by the InternVL2-style sampling method [43]. Specifically, each image is first scaled and padded to the closest HD resolution such that both height and width are divisible by 336. The padded image is then divided into non-overlapping crops of size 336×336, with a maximum of 16 crops per image. In addition to these local crops, we include a globally resized version of the full image (also 336×336) to capture holistic context. Each of these crops, including the resized full image, is independently passed through the image encoder, yielding 577 tokens (based on $14\times 14$ patch sizes) per crop, including the [CLS] token, each with a hidden dimension of 1024. We extract the [CLS] token from each encoded crop and compute their mean across all crops and the full image. This averaged [CLS] representation is then used as input to the classification head.
10.2 Implementation Details
@@
We performed all experiments in PyTorch [44] using HuggingFace Transformers [58] library (v4.41.2). Training employs the AdamW [30] optimizer with a constant learning rate of $1\times 10^{-5}$, and cross-entropy loss. We used a batch size of 256 for variants without dynamic cropping, and 64 for those with dynamic cropping. To improve stability, gradient clipping with a maximum norm of 1.0 was applied. Each model is trained for up to 45 epochs with early stopping based on validation accuracy. We use a lightweight classification head (two fully connected layers with non-linear activations) to predict one of twelve rotation classes. All training was performed on a single NVIDIA H100 GPU, which has an average latency of 0.5s per image during inference. The same model has a latency of around 0.9s on a single NVIDIA A100 GPU. We used the implementation of OCRBenchV2333https://github.com/Yuliang-Liu/MultimodalOCR/blob/main/OCRBench_v2/eval_scripts/vqa_metric.py to compute the ANLS metrics. We adapted this script to report both word-level and character-level ANLS metrics. For structured analysis on ORB-SROIE, we test the model against 4 fields: company, date, address and total for the percentage match calculations. We provide more details about the configurations used across different baselines.
Tesseract [49]
@@
LLaMA-4 (Maverick) 17B [54]
@@
Gemini-2.5 Flash [16]
For the Gemini-2.5 Flash baseline, we used Google’s multimodal model accessed via the google.genai Python SDK. The model gemini-2.5-flash was configured with an authenticated API key and executed in streaming mode for OCR text extraction. Each image was read as binary data and sent to the model as an inline MIME-typed input. The outputs were returned as plain-text responses and stored incrementally in JSONL format for further evaluation.
@@
  • Access: google.genai SDK
@@
10.3 Data Sources and Augmentation
For English training data, we use documents from various categories such as invoices, rent agreements, electricity bills, and gas bills, collected from internal repositories. For Indic data, we curate document images from 11 languages (Hindi, Bengali, Tamil, Malayalam, Telugu, Marathi, Kannada, Gujarati, Punjabi, Oriya, and Assamese) using Wikisource. To ensure fair evaluation, we strictly exclude all test images from the training sets. Specifically, the ORB-En and ORB-Indic benchmarks are fully disjoint from the training data. Moreover, we do not use the SROIE 2019 or the FUNSD training split during model training. All results reported are therefore in a zero-shot setting, highlighting the generalization capability of the proposed method across both English and Indic scripts on our benchmark.
@@ 10.4 Evaluation and Baselines
    ORB-En: A 897-image benchmark composed of 347 images from the SROIE 2019 test set, 500 images from the SynthDog validation set and 50 images from the FUNSD validation set, uniformly rotated and re-annotated.
@@
    ORB-Indic: A 966-image benchmark covering 11 Indic scripts across 12 different angles, each manually annotated and verified for rotation labels, along with corresponding ground-truth transcriptions for OCR evaluation.
@@
  • Tesseract [49]
  • TrOCR [33]
  • docTR [34]
  • Nanonets-OCR-s [41]
  • Gemma-3 27B [52]
  • LLaMA-4 (Maverick) 17B [54]
  • GPT-4o [24]
@@
  • Gemini-2.5 Pro [17]
  • Gemini-2.5 Flash [16]
The English model, trained exclusively on Latin-script documents does not transfer well to Indic scripts, achieving only 77.4% accuracy on the ORB-Indic benchmark.
10.5 Pipeline Integration
@@
To measure downstream impact, we insert our rotation classifier before OCR engines. We compare OCR related metrics with and without rotation correction. Field-level accuracy is used for structured forms like ORB-En (SROIE), and string alignment metrics are used for the free-text Indic dataset ORB-Indic, ORB-En (SynthDog) and ORB-En (FUNSD) data.
Model ORB-En ORB-Indic
TrOCR 26.22 68.90
docTR 63.28 46.95
Nanonets-OCR-s 24.83 37.50
Gemma-3 27B 30.57 33.23
LLaMA-4 (Maverick) 17B 65.49 56.70
GPT-4o 59.58 50.60
Gemini-2.5 Flash 39.55 41.46
Gemini-2.5 Pro 34.11 40.85
Ours 96.81 92.68
Table 7: Accuracy (%) for 4-class document rotation classification on ORB-En and ORB-Indic. Best in bold, second-best underlined.
11 Additional Ablation Study
Additionally, we also conducted a second ablation study of the effect of the number of patches on the accuracy of rotation-class classification of the 12 class ORB-En FunSD dataset. This shows that while the accuracy increases as we increase the number of patches, it plateaus at around 32 patches, which has a latency of around 1.8 times the base dynamic cropping method, while yielding the same accuracy percentage, which makes the use of dynamic cropping a better option rather than specifying the patch size to be used explicitly.
Patch count Accuracy (%) $\uparrow$
Single crop (1 patch) 60
Moderate cropping (8 patches) 86
Dense cropping (32 patches) 98
Dynamic cropping (Ours) 98
Table 8: Ablation study on the number of dynamic cropping patches for 12-class document rotation classification on the ORB-En FUNSD dataset.
Model Ideal upright conditions w/o Rotation w/ Rotation
ORB-En-SynthDog dataset — 4 class rotation
docTR 56.23 40.04 88.03 69.07 56.23 40.04
Nanonets-OCR-s 2.61 1.73 44.14 34.25 2.61 1.73
Gemma-3 27B 61.56 30.29 75.15 50.06 61.56 30.29
LLaMA-4 (Maverick) 17B 29.37 14.09 59.91 41.75 29.37 14.09
Qwen-2.5-VL 32.20 4.67 55.10 33.65 34.98 7.77
mPLUG-DocOwl2 85.45 80.87 93.35 87.55 86.44 81.82
GPT-4o 82.22 73.65 92.35 82.08 82.22 73.65
Gemini-2.5 Pro 58.58 52.34 70.96 65.41 67.07 61.16
Gemini-2.5 Flash 22.43 15.33 26.24 18.66 22.43 15.33
ORB-Indic dataset — 4 class rotation
Nanonets-OCR-s 75.78 62.12 92.07 87.28 75.78 62.12
Gemma-3 27B 75.43 48.19 92.01 75.53 76.16 49.41
LLaMA-4 (Maverick) 17B 47.25 24.73 84.41 68.83 47.25 24.73
GPT-4o 82.43 63.76 94.55 89.14 82.00 64.14
Chitrapathak 34.56 16.10 82.33 68.51 34.56 16.10
Gemini-2.5 Pro 48.48 39.10 67.50 58.76 56.36 47.33
Gemini-2.5 Flash 26.33 11.93 27.17 12.80 27.16 12.84
Table 9: ANLS (Average Normalized Levenshtein Similarity) under 4-class rotation across English and Indic datasets for different models. Word- and character-level scores are reported for ideal upright inputs, OCR without rotation correction, and OCR with rotation correction (lower is better).
@@ 11 Additional Ablation Study
Table 10: Performance comparison on ORB-En (SynthDog) and 4-class ORB-Indic documents with and without the proposed rotation correction pipeline, alongside ideal conditions. We report Word Error Rate (WER) and Character Error Rate (CER) on free-form outputs from both open-source and closed-source models (lower is better), computed using the HuggingFace evaluate function. Second best scores are underlined.
@@
Table 11: Performance comparison on ORB-En (SynthDog) and 4-class ORB-Indic documents with and without the proposed rotation correction pipeline, alongside ideal conditions. We report percentage match scores at word and character level (higher is better). Second best scores are underlined.